# On the Theoretical Foundation of Sparse Dictionary Learning in Mechanistic Interpretability

### v2

@@ 3.2 Polysemanticity and Local Minima
Under this condition, neuron 1 activates for both features while neuron 2 remains dead. Let $a_{1}=\langle w_{e}^{1},w_{p}^{1}\rangle$ and $a_{2}=\langle w_{e}^{1},w_{p}^{2}\rangle$. The optimal decoder is:
$w_{d}^{1,*}=\frac{M_{1}a_{1}w_{r}^{1}+M_{2}a_{2}w_{r}^{2}}{M_{1}a_{1}^{2}+M_{2}a_{2}^{2}}$ (19)
Since $w_{r}^{1}\not\propto w_{r}^{2}$, the resulting loss is:
$\tilde{\mathcal{L}}_{\text{SDL}}^{*}=\frac{M_{1}M_{2}(M_{2}a_{2}^{2}+M_{1}a_{1}^{2})\|a_{2}w_{r}^{1}-a_{1}w_{r}^{2}\|^{2}}{(M_{1}a_{1}^{2}+M_{2}a_{2}^{2})^{2}}>0$ (20)
The spurious local minimum in Example 3.3 demonstrates a prevalent failure mode for SDL methods: a single SDL latent neuron becomes polysemantic, responding to multiple ground-truth features, while other neurons remain dead. This occurs when optimization becomes trapped in a local minimum where gradient descent cannot escape. We now formalize and characterize this polysemanticity.
Definition 3.4 (Polysemanticity Pattern).
@@
A polysemanticity pattern is a collection $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ where each $\mathcal{F}_{i}\subseteq[n]$ represents a subset of ground-truth features. The pattern is called polysemantic if $\exists i,|\mathcal{F}_{i}|\geq 2$.
Given an SDL configuration with activation function $\sigma$ and projection vectors $\{w_{p}^{d}\}_{d=1}^{n}$, a polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forming a partition of $[n]$ is called realizable if there exists an encoder $W_{e}=[(w_{e}^{1})^{\top},\ldots,(w_{e}^{n_{q}})^{\top}]^{\top}$ such that:
$\forall i\in[n_{q}],\quad\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{e}w_{p}^{d})\right)_{i}>0\right\}$ (21)
@@
That is, neuron $i$ activates for exactly the features in $\mathcal{F}_{i}$ when they are active in isolation.
Theorem 3.5 (Existence of Spurious Local Minima).
Consider $\tilde{\mathcal{L}}_{\text{SDL}}(W_{d},W_{e})=\sum_{d=1}^{n}M_{d}\left\|w_{r}^{d}-W_{d}\sigma(W_{e}w_{p}^{d})\right\|^{2}$ under Assumptions 2.3 and 2.4 with $n\geq 2$ and $n_{q}\geq n$.
For any realizable polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ with at least one $|\mathcal{F}_{i}|\geq 2$, there exists a local minimum point $(W_{d},W_{e})$ of the framework exhibiting this pattern.
3.3 Feature Absorption and Hierarchical Concepts
Empirical observations suggest that feature absorption is prevalent during the training of SDL (Chanin et al., 2025). We argue that this phenomenon arises when ground-truth features exhibit hierarchical structure and can be considered as a special type of realizable polysemanticity pattern.
@@ 3.3 Feature Absorption and Hierarchical Concepts
Consider a representation space containing animal concepts. Ground-truth features include a parent concept ”Dog” and four sub-concepts: ”Border Collie”, ”Golden Retriever”, ”Husky”, and ”German Shepherd”.
As illustrated in Figure 5, in the ideal case, SDL would learn four separate monosemantic neurons for Dog, Cat, Horse, and Elephant, with each dog breed activating only the Dog neuron. However, feature absorption can occur during training, resulting in the pattern shown in Figure 5 (right): one neuron exclusively captures ”Border Collie” (the absorbed feature), while a second neuron responds to the remaining three breeds collectively (the main line interpretation of ”Dog”).
@@
A set of ground-truth features exhibits hierarchical structure if it contains a $p$ and a set of sub-concepts $c_{1},\ldots,c_{k}$ satisfying: $p(x)>0\iff\exists i\in[k],c_{i}(x)>0$.
Theorem 3.8 (Existence of Feature Absorption).
Any ground truth features exhibiting hierarchical structure correspond to a polysemanticity pattern with feature absorption. If realizable, there is a local minimum point exhibiting it.
Theorem 3.5 reveals the prevalence of local minima in SDL training that challenge gradient-based optimization and their connection to polysemantic neurons. Theorem 3.8 shows that hierarchical concept structures naturally give rise to feature absorption patterns, which manifest as local minima when realizable, explaining its empirical prevalence.
4 The Feature Anchoring
Motivated by our theoretical results, we propose Anchored Sparse Autoencoders (See Figure 6) to guide optimization toward the global optimum. Specifically, we constrain the first $k$ encoder rows and decoder columns to match ground-truth feature directions via an auxiliary loss to mitigate the undetermined nature of the SDL optimization problem:
@@ 4 The Feature Anchoring
$\begin{split}\mathcal{L}_{\text{anchor}}&=\|\mathbf{W}_{e}[1{:}k,:]-\tilde{\mathbf{W}}_{p}[:,1{:}k]^{\top}\|^{2}_{F}\\ &+\|\mathbf{W}_{d}[:,1{:}k]-\tilde{\mathbf{W}}_{r}[:,1{:}k]\|^{2}_{F}\end{split}$ (22)
These anchor features are obtained either from ground-truth when available (Linear Representation Bench) or by computing mean embeddings of dataset subpopulations (Luo et al., 2024). During the evaluation, we only compute the recovery metric for features that were not selected as anchors.
Figure 6: Anchored SAE. First $k$ neurons (highlighted) constrained to ground-truth via $\mathcal{L}_{\text{anchor}}$; other neurons learn via SDL.
@@ 5 Experimental Results
5.1 The Linear Representation Bench
To validate our theoretical predictions, we design the Linear Representation Bench, a benchmark that precisely instantiates Assumptions 2.3 and 2.4. We generate synthetic representations $\mathbf{x}_{p}(s)=\mathbf{W}_{p}\mathbf{x}(s)$ where ground-truth features $\mathbf{x}(s)\in\mathbb{R}^{n}$ follow shifted exponential distributions with sparsity $\mathcal{S}$. The feature matrix $\mathbf{W}_{p}\in\mathbb{R}^{n_{p}\times n}$ is constructed via gradient-based optimization to satisfy bounded interference $|\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{p}\rangle|\leq\epsilon$ or negative interference $\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{p}\rangle\leq 0$ with unit-norm columns. We evaluate SDL methods by measuring feature recovery: $\mathbb{E}_{i}[\max_{j}|\langle\mathbf{w}^{i}_{p},\mathbf{w}^{j}_{e}\rangle|]$, the expected maximum alignment between each ground-truth feature $\mathbf{w}^{i}_{p}$ and learned encoder columns $\mathbf{w}^{j}_{e}$. As shown in Table 1, our proposed Anchored SAE significantly improves feature recovery performance.
Interference n Method
SAE Anchored SAE
Bounded 1000 0.256 0.284
Negative 1000 0.235 0.260
Table 1: We compare standard SAE and Anchored SAE (with $k=30$ anchor features) on synthetic representations in $d=768$ dimensions with $n\in\{1000,1500,2000\}$ features (Appendix C).
5.2 Neuron Resampling Helps Escape Local Minima
Our local minima analysis (Theorem 3.5) provides a novel explanation for why dead neuron resampling (Bricken et al., 2023) improves SDL training. Dead neurons ($\mathcal{F}_{i}=\emptyset$) contribute zero gradient in spurious local minima, leaving optimization trapped. Resampling reinitializes these neurons toward under-reconstructed directions, perturbing the optimization away from the local basin. To validate this, we conduct experiments on the Llama 3 model (Figure 7). Resampling enables the optimizer to escape spurious local minima, achieving lower reconstruction loss.
@@ 5.2 Neuron Resampling Helps Escape Local Minima
Figure 7: Feature resampling accelerates convergence and improves final loss. Training curves on Llama 3 comparing standard SAE training (blue) with periodic dead neuron resampling (red).
6 Related Works
6.1 Mechanistic Interpretability

### v3

@@ 3.4 Characterizing Spurious Partial Minima of SDL
Consider the configuration where neuron 1 activates for both features while neuron 2 remains dead:
$W_{E}^{*}=\begin{bmatrix}1&1\\ 0&0\end{bmatrix},\quad W_{D}^{*}=\begin{bmatrix}1/2&0\\ 1/2&0\end{bmatrix}$ (18)
Within the activation pattern region $\Omega_{\mathcal{A}}=\{W_{E}:\langle w_{E}^{1},w_{p}^{1}\rangle>0,\langle w_{E}^{1},w_{p}^{2}\rangle>0\}$, direct calculation shows $\nabla_{W_{D}}\tilde{\mathcal{L}}=0$ and $\nabla_{W_{E}}\tilde{\mathcal{L}}=0$. By traditional biconvex optimization theory (Gorski et al., 2007), $(W_{D}^{*},W_{E}^{*})$ is a partial optimum. However, this configuration exhibits polysemanticity with suboptimal loss: $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=1>0$.
Definition 3.6 (Activation Pattern).
@@
An activation pattern is a collection $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{N})$ where $\mathcal{F}_{i}\subseteq[n]$ denotes the set of ground-truth features that activate neuron $i$.
An activation pattern $\mathcal{P}$ is called:
  • Polysemantic if $\exists i\in[N]$ such that $|\mathcal{F}_{i}|\geq 2$ (at least one neuron responds to multiple features).
  • Realizable if there exists an encoder $\sigma(W_{E}(\cdot))\in\mathbb{R}^{N\times n_{p}}$ such that:
    $\forall i\in[N],\quad\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{E}w_{p}^{d})\right)_{i}>0\right\}$ (19)
@@
    That is, neuron $i$ activates for exactly the features in $\mathcal{F}_{i}$ when they appear in isolation.
Theorem 3.7 (Prevalence of Spurious Partial Minima).
Under Assumptions 2.3 and 2.4 with $n\geq 2$ and $n_{q}\geq n$, for any activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ that is realizable, polysemantic, and forms a partition of $[n]$, there exists a partial minimum $(W_{D}^{*},W_{E}^{*})$ of $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibiting this pattern with positive loss.
This establishes that partial minima are pervasive in SDL optimization: every realizable polysemantic activation pattern corresponds to a partial optima point where gradient descent can become trapped, a persistent challenge for SDL.
3.5 Theoretical Explanation for Feature Absorption
Feature absorption—where one neuron captures, or ”absorbs”, a specific sub-concept while another responds to remaining related concepts—frequently occurs in SDL training (Chanin et al., 2025). Though prevalently encountered and unwanted, researchers have limited understanding about why feature absorption occurs in SDL training. Here we show hierarchical concept structures naturally introduce realizable activation patterns exhibiting feature absorption and therefore connected with the framework’s partial minima.
@@ 3.5 Theoretical Explanation for Feature Absorption
Consider a representation space with a parent concept ”Dog” and four sub-concepts: ”Border Collie”, ”Golden Retriever”, ”Husky”, and ”German Shepherd”. Ideally, SDL learns separate monosemantic neurons for Dog, Cat, Horse, and Elephant, with each dog breed activating only the Dog neuron. However, feature absorption can result in the pattern in Figure 5: one neuron exclusively captures ”Border Collie” (absorbed feature), while another responds to the remaining three breeds collectively (main line interpretation).
@@
A set of ground-truth features exhibits hierarchical structure if it is composed of a parent concept $p$ and a set of sub-concepts $c_{1},\ldots,c_{k}$ satisfying: $p(x)>0\iff\exists i\in[k],c_{i}(x)>0$.
Theorem 3.10 (Feature Absorption from Hierarchical Structure).
Suppose there exist $M$ parent concepts with hierarchical decompositions into sub-concepts: for each $i\in[M]$, parent concept $d_{i}$ decomposes into sub-concepts $\mathcal{F}_{i}=\{d_{i,1},\ldots,d_{i,k_{i}}\}$ where $k_{i}\geq 2$.
If the activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{M})$ is realizable, then $\forall i^{*}\in[M]$, $\exists j^{*}\in[k_{i^{*}}]$ such that the following activation pattern exhibiting feature absorption is realizable:
$\mathcal{P}^{\prime}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{i^{*}}\setminus\{d_{i^{*},j^{*}}\},\{d_{i^{*},j^{*}}\},\ldots,\mathcal{F}_{M})$ (20)
Theorem 3.10 explains why feature absorption persists even with careful training: hierarchical concept structures naturally induce realizable polysemantic patterns that manifest as partial minima, providing theoretical grounding for this widely observed empirical phenomenon.
4 Method
Our theoretical analysis reveals a fundamental challenge in SDL optimization: the underdetermined nature of the loss landscape (Theorem 3.4), and the solution space admits multiple configurations—some achieving zero reconstruction loss without recovering any interpretable ground-truth features (Figure 4). These theoretical findings motivate our method, feature anchoring, a technique that constrains a subset of features to align with known semantic directions.
4.1 Anchor Feature Extraction
Feature anchoring requires identifying $k$ anchor features $\{\tilde{\mathbf{w}}_{p}^{(i)},\tilde{\mathbf{w}}_{r}^{(i)}\}_{i=1}^{k}$ that represent semantically meaningful directions in the representation space. We present two methods for obtaining these anchors:
Ground-Truth Features (Linear Representation Bench).
When ground-truth features are available—as in our Linear Representation Bench where features are known by construction—we directly use a random subset of $k$ ground-truth feature directions:
$\tilde{w}_{p}^{(i)}=W_{p}^{\text{true}}[:,i],\quad\tilde{w}_{r}^{(i)}=W_{r}^{\text{true}}[:,i],\quad i\in\mathcal{K}$ (21)
where $\mathcal{K}\subset[n]$ is a randomly selected subset of size $k$.
Subpopulation Mean Embeddings.
For real-world datasets where ground-truth features are unknown, we identify semantic subpopulations Luo et al. (2024) and compute their mean representations. Specifically, given a labeled dataset $\mathcal{D}=\{(s_{j},y_{j})\}_{j=1}^{N}$, where $y_{j}\in\{1,\ldots,C\}$, we compute the mean of representations for each class:
$\bar{\mathbf{x}}_{p}^{(c)}=\frac{1}{|\{j:y_{j}=c\}|}\sum_{j:y_{j}=c}\mathbf{x}_{p}(s_{j})$ (22)
Then we normalize each mean representation to obtain anchor directions:
$\tilde{\mathbf{w}}_{p}^{(c)}=\frac{\bar{\mathbf{x}}_{p}^{(c)}}{\|\bar{\mathbf{x}}_{p}^{(c)}\|_{2}}$ (23)
4.2 SDL Loss Function with Feature Anchoring
Given $k$ anchor features $\{\tilde{\mathbf{w}}_{p}^{(i)},\tilde{\mathbf{w}}_{r}^{(i)}\}_{i=1}^{k}$, we modify the SDL optimization objective to include an anchoring penalty that constrains the first $k$ encoder rows and decoder columns to align with these anchors.
The complete anchored SDL objective is:
$\mathcal{L}_{\text{SDL-FA}}=\mathcal{L}_{\text{SDL}}+\lambda_{\text{anchor}}\mathcal{L}_{\text{anchor}}$ (24)
where $\mathcal{L}_{\text{SDL}}$ is the standard SDL loss (Equation 3) and the anchoring loss is:
@@ 4.2 SDL Loss Function with Feature Anchoring
$\begin{split}\mathcal{L}_{\text{anchor}}&=\|W_{E}[1{:}k,:]-[\tilde{w}_{p}^{(1)},\ldots,\tilde{w}_{p}^{(k)}]^{\top}\|^{2}_{F}\\ &\quad+\|W_{D}[:,1{:}k]-[\tilde{w}_{r}^{(1)},\ldots,\tilde{w}_{r}^{(k)}]\|^{2}_{F}\end{split}$ (25)
This feature anchoring technique (Equation 25) reduces the underdetermined nature of SDL training, and is method-agnostic: it applies equally to SAEs, transcoders, crosscoders, and their variants (TopK SAEs, Matryoshka SAEs, etc.) since it only constrains the encoder $W_{E}$ and decoder $W_{D}$ matrices that are common to all SDL architectures. This universality makes feature anchoring a broadly useful technique for improving performance across SDL methods.
@@ 5 Experimental Results
We first evaluate feature anchoring with various SDL methods on the Linear Representation Bench (Section 5.1), a synthetic benchmark with fully accessible ground-truth features that precisely instantiates our theoretical assumptions. Second, we validate feature anchoring on real-world data by training SDL methods on CLIP embeddings of ImageNet-1K. Third, we provide empirical evidence showing that dead neuron resampling helps escape spurious local minima in large language models. We provide comprehensive ablation studies in Appendix G.
5.1 Results on The Linear Representation Bench
To validate our theoretical predictions, we design the Linear Representation Bench, a benchmark that precisely instantiates Assumptions 2.3 and 2.4 with fully known GT features.
Data Generation. We generate synthetic representations $\mathbf{x}_{p}(s)=W_{p}^{\text{true}}\mathbf{x}(s)$ where ground-truth features $\mathbf{x}(s)\in\mathbb{R}^{n}$ follow shifted exponential distributions with sparsity $\mathcal{S}$. The feature matrix $W_{p}^{\text{true}}\in\mathbb{R}^{n_{p}\times n}$ is constructed via gradient-based optimization (Detailed in Appendix C).
Metrics. Given learned features $W_{p}^{\text{learned}}=W_{E}^{\top}\in\mathbb{R}^{n_{p}\times n_{q}}$ and ground truth features $W_{p}^{\text{true}}\in\mathbb{R}^{n_{p}\times n}$ (columns are unit-norm feature directions), we compute the similarity matrix $\mathbf{S}=|W_{p}^{\text{learned}\top}W_{p}^{\text{true}}|\in\mathbb{R}^{n_{q}\times n}$. For each ground truth feature $i\in[n]$, we find its best match: $s_{i}=\max_{j}S_{ji}$.
  • GT Recovery is the fraction of ground truth features with $s_{i}>\tau$, $\mathbf{M_{GT}}(W_{p}^{\text{learned}})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{s_{i}>\tau\}$.
  • Maximum Inner Product is the mean of best matches: $\mathbf{M_{IP}}(W_{p}^{learned})=\frac{1}{n}\sum_{i=1}^{n}s_{i}$.
As shown in Table 1, feature anchoring significantly improves feature recovery performance for all SDL methods.
Table 1: Feature recovery on the Linear Representation Bench ($n=1000$, $n_{p}=n_{r}=768$, $n_{q}=16000$, $\mathcal{S}=0.99$). Feature anchoring (FA) consistently improves GT Recovery and Maximum Inner Product (MIP) across all SDL methods.
[…]
ReLU SAE 0.00% 0.205 2787.4
   + Feature Anchoring 0.00% 0.246 1715.1
JumpReLU SAE 0.00% 0.237 1572.3
   + Feature Anchoring 0.00% 0.333 1139.6
TopK SAE 84.90% 0.983 54.1
   + Feature Anchoring 87.63% 0.986 54.1
BatchTopK SAE 84.80% 0.981 65.0
   + Feature Anchoring 89.38% 0.988 60.1
Matryoshka SAE 83.70% 0.982 58.9
   + Feature Anchoring 87.32% 0.985 57.9
Transcoder 23.60% 0.838 50.6
   + Feature Anchoring 25.05% 0.838 49.5
Crosscoder 56.42% 0.940 83.4
   + Feature Anchoring 57.71% 0.941 116.3
5.2 Results on CLIP Embeddings of ImageNet
To validate that feature anchoring generalizes beyond synthetic benchmarks to real-world representations, we apply our method to CLIP embeddings of ImageNet-1K (Russakovsky et al., 2015). We extract CLIP-ViT-B/32 (Radford et al., 2021) image embeddings for all ImageNet training images and compute ground-truth anchor features as normalized class mean embeddings: $\tilde{w}_{p}^{(c)}=\bar{x}_{p}^{(c)}/\|\bar{x}_{p}^{(c)}\|_{2}$ where $\bar{x}_{p}^{(c)}$ is the mean embedding over all images in class $c$. We train TopK, BatchTopK, and Matryoshka SAEs with $n_{p}=n_{r}=768$, $n_{q}=16,384$, and use $k=30$ anchors with $\lambda_{\text{anchor}}=1.0$. Table 2 shows feature anchoring consistently improves feature recovery (Examples in Appendix H).
Table 2: Feature recovery on the CLIP embeddings.
[…]
TopK SAE 0.00% 0.517 64.0
   + Feature Anchoring 24.13% 0.851 64.0
BatchTopK SAE 0.00% 0.508 64.0
   + Feature Anchoring 24.13% 0.847 64.0
Matryoshka SAE 0.00% 0.683 128.0
   + Feature Anchoring 24.45% 0.858 128.0
5.3 Neuron Resampling Helps Escape Partial Minima
Our partial minima analysis (Theorem 3.7) proves the prevalence of spurious partial minima and connects them to dead neurons ($\mathcal{F}_{i}=\emptyset$). To address this, we utilize neuron resampling (Bricken et al., 2023) to help SDL training overcome these partial minima. We argue that resampling reinitializes dead neurons toward under-reconstructed directions, perturbing the optimization away from the partial minima.
To validate this, we train SAEs on Llama 3.1 8B Instruct (layer 12, dimension 4096) with latent dimension 131,072 for 30,000 steps on FineWeb-Edu (Penedo et al., 2024), processing 4096 token activations per step. We compare standard training against feature resampling after 5000 steps. As shown in Figure 6, resampling enables the optimizer to escape spurious local minima, achieving lower final loss.
@@ 5.3 Neuron Resampling Helps Escape Partial Minima
Figure 6: Feature resampling accelerates convergence and improves final loss. Training curves on Llama 3.1 8B comparing standard SAE training (blue) with periodic dead neuron resampling after 5000 steps (in the plot x axis is scaled by 100).
6 Conclusion
We develop the first unified theoretical framework for Sparse Dictionary Learning in mechanistic interpretability, demonstrating how diverse SDL methods instantiate a single optimization problem. We prove that SDL exhibits piecewise biconvex structure, bridging mechanistic interpretability with classical optimization theory. We characterize the global minimum and establish that the optimization is fundamentally underdetermined, admitting solutions that achieve zero reconstruction loss without recovering interpretable features. We demonstrate that spurious partial minima exhibiting polysemanticity are pervasive, and prove that hierarchical concept structures naturally induce feature absorption patterns that manifest as partial minima. To validate our theory, we design the Linear Representation Bench with fully accessible ground-truth features, and propose feature anchoring—a technique applicable to all SDL methods that addresses the underdetermined nature of optimization.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Code Availability
We provide full scripts of the Linear Representation Bench and feature anchoring in this GitHub repository: The_Linear_Representation_Bench.
Appendix A Taxonomy for Sparse Dictionary Learning in Mechanistic Interpretability
Figure 7: Hierarchical taxonomy of Sparse Dictionary Learning research in Mechanistic Interpretability.
Appendix B Notations
We summarize the key notations used throughout this paper.
Notation Description
Distributions and Indexing
$\mathcal{D}$ Distribution over inputs
$\mathcal{X}$ Input space
$s$ Sample drawn from $\mathcal{D}$
$\mathbf{e}_{d}$ Standard basis vector (1 in position $d$, 0 elsewhere)
Representations and Dimensions
$\mathbf{x}(s)$ Ground-truth features for input $s\sim\mathcal{D}$
$\mathbf{x}_{p}(s)$ Input representation to SDL model (observed)
$\mathbf{x}_{r}(s)$ Target/output representation for SDL model
$\mathbf{x}_{q}(s)$ Latent activations in SDL bottleneck
$n$ Number of ground-truth features (dimension of $\mathbf{x}$)
$n_{p}$ Dimension of input representation $\mathbf{x}_{p}$
$n_{r}$ Dimension of target representation $\mathbf{x}_{r}$
$n_{q}$ Dimension of latent space $\mathbf{x}_{q}$
SDL Architecture
$W_{E}$ Encoder matrix ($n_{q}\times n_{p}$)
$W_{D}$ Decoder matrix ($n_{r}\times n_{q}$)
$\sigma(\cdot)$ Sparsity-inducing activation function
$\mathbf{w}_{E}^{i}$ $i$-th row of encoder $W_{E}$ (encodes to neuron $i$)
$\mathbf{w}_{D}^{i}$ $i$-th column of decoder $W_{D}$ (decodes from neuron $i$)
Linear Representation Hypothesis
$W_{p}$ Feature matrix for $\mathbf{x}_{p}$ ($n_{p}\times n$)
$W_{r}$ Feature matrix for $\mathbf{x}_{r}$ ($n_{r}\times n$)
$\mathbf{w}_{p}^{d}$ $d$-th column of $W_{p}$ (feature direction for feature $d$)
$\mathbf{w}_{r}^{d}$ $d$-th column of $W_{r}$ (feature direction for feature $d$)
$W_{p}^{\text{true}}$ Ground-truth feature matrix (Linear Representation Bench)
$W_{p}^{\text{learned}}$ Learned feature matrix ($W_{E}^{\top}$)
Sparsity and Interference
$S$ Sparsity level: $\Pr(\mathbf{x}_{i}(s)=0)\geq S$
$M$ Maximum interference: $\max_{i\neq j}\langle W_{p}[:,i],W_{p}[:,j]\rangle$
$M_{d}$ Weight for feature $d$: $\Pr(\mathbf{x}(s)=x_{d}(s)\mathbf{e}_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$
Feature Anchoring
$k$ Number of anchor features
$\tilde{\mathbf{w}}_{p}^{(i)},\tilde{\mathbf{w}}_{r}^{(i)}$ $i$-th anchor feature pair
$\lambda_{\text{anchor}}$ Anchoring loss weight
$\mathcal{K}$ Set of indices for selected anchor features
Loss Functions
$\mathcal{L}_{\text{SDL}}$ Standard SDL reconstruction loss
$\tilde{\mathcal{L}}_{\text{SDL}}$ Approximate SDL loss (extreme sparsity regime)
$\mathcal{L}_{\text{anchor}}$ Feature anchoring penalty
$\mathcal{L}_{\text{SDL-FA}}$ Anchored SDL objective: $\mathcal{L}_{\text{SDL}}+\lambda_{\text{anchor}}\mathcal{L}_{\text{anchor}}$
Activation Patterns
$\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{N})$ Activation pattern
$\mathcal{F}_{i}$ Set of ground-truth features activating neuron $i$
$\mathcal{A}(d)$ Set of neurons activated by feature $d$
$\Omega_{\mathcal{A}}$ Activation pattern region (fixed activation structure)
$c$ Activation threshold for $\sigma$
Evaluation Metrics
$\mathbf{M}_{\text{GT}}$ GT Recovery: fraction of features with $\tau$ alignment
$\mathbf{M}_{\text{IP}}$ Maximum Inner Product: mean best alignment per feature
$\tau$ Threshold for GT Recovery (typically $\tau=0.9$ or $0.95$)
$\mathbf{S}$ Similarity matrix: $|W_{p}^{\text{learned}\top}W_{p}^{\text{true}}|$
Appendix C The Linear Representation Bench
The Linear Representation Bench is a synthetic benchmark designed to precisely instantiate the Linear Representation Hypothesis (Assumption 2.3) and Representation Assumptions (Assumption 2.4) with fully accessible ground-truth features (See Figure 8 for a visualization of the benchmark). This enables rigorous evaluation of SDL methods under controlled conditions where feature recovery can be directly measured.
C.1 Ground-Truth Feature Matrix Generation
We construct the feature matrix $\mathbf{W}_{p}\in\mathbb{R}^{n_{p}\times n}$ to satisfy the unit-norm and bounded interference conditions (Assumption 2.4).
Initialization.
Initialize $\mathbf{W}_{p}$ with random Gaussian entries and normalize each column to unit $\ell_{2}$-norm:
$\mathbf{w}_{p}^{d}\leftarrow\frac{\mathbf{w}_{p}^{d}}{\|\mathbf{w}_{p}^{d}\|_{2}},\quad\forall d\in[n]$ (26)
Interference Minimization.
We minimize pairwise interference via projected gradient descent. Define the soft-thresholded interference loss:
$\mathcal{L}_{\text{int}}(\mathbf{W}_{p})=\sum_{i\neq j}\left[\max\left(0,\langle\mathbf{w}_{p}^{i},\mathbf{w}_{p}^{j}\rangle-(M-\epsilon)\right)\right]^{2}+\lambda\sum_{i\neq j}\left[\max\left(0,\langle\mathbf{w}_{p}^{i},\mathbf{w}_{p}^{j}\rangle\right)\right]^{2}$ (27)
where $\epsilon>0$ is a tolerance margin and $\lambda>0$ is a small regularization weight encouraging negative interference.
At each iteration, we compute the gradient with respect to $\mathbf{W}_{p}$, perform a gradient step, and project back to the unit sphere:
$\mathbf{W}_{p}\leftarrow\mathbf{W}_{p}-\eta\nabla_{\mathbf{W}_{p}}\mathcal{L}_{\text{int}},\quad\mathbf{w}_{p}^{d}\leftarrow\frac{\mathbf{w}_{p}^{d}}{\|\mathbf{w}_{p}^{d}\|_{2}},\quad\forall d\in[n]$ (28)
This procedure continues until the maximum interference satisfies $\max_{i\neq j}\langle\mathbf{w}_{p}^{i},\mathbf{w}_{p}^{j}\rangle\leq M$.
C.2 Sparse Coefficient Generation
For each sample $i\in[N]$, we generate sparse ground-truth features $\mathbf{x}^{(i)}\in\mathbb{R}^{n}_{+}$ as follows:
Sparsity Mask.
Each feature $d\in[n]$ is independently activated with probability $(1-\mathcal{S})$:
$m_{d}^{(i)}\sim\text{Bernoulli}(1-\mathcal{S})$ (29)
Feature Magnitudes.
Active features follow a shifted exponential distribution:
$x_{d}^{(i)}=m_{d}^{(i)}\cdot\left(c_{\min}+\text{Exp}(\beta)\right)$ (30)
where $c_{\min}\geq 0$ is a minimum activation threshold and $\beta>0$ is the scale parameter.
This construction ensures:
  • Non-negativity: $x_{d}^{(i)}\geq 0$ for all $d,i$
  • Sparsity: $\Pr(x_{d}^{(i)}=0)=\mathcal{S}$ for all $d$
  • Independence: Features $\{x_{d}^{(i)}\}_{d=1}^{n}$ are mutually independent
C.3 Data Synthesis
The observed representations are computed via linear combination:
$\mathbf{x}_{p}^{(i)}=\mathbf{W}_{p}\mathbf{x}^{(i)}=\sum_{d=1}^{n}x_{d}^{(i)}\mathbf{w}_{p}^{d}$ (31)
By construction, this dataset exactly satisfies Assumptions 2.3 and 2.4, enabling evaluation of SDL methods with complete knowledge of ground-truth features.
C.4 Default Configuration
Table 3 lists the default parameters used in our experiments.
Table 3: Default configuration for the Linear Representation Bench.
Parameter Symbol Default Value
Number of features $n$ 1000
Representation dimension $n_{p}$ 768
[…]
Number of samples NN 100,000
Sparsity level $\mathcal{S}$ 0.99
Maximum interference $M$ 0.1
This controlled setting enables direct measurement of feature recovery metrics with ground-truth access, complementing evaluations on real neural network representations where true features are unknown.
Appendix D Related Works
D.1 Mechanistic Interpretability